### v1

@@ 4.2 Determining stable nodes
In this section, we work with stack-free phylogenetic networks, denoted by $\mathcal{N}=(V,E)$. Proposition 2 shows that if two reticulation-visible networks have equal $\mu$-representations, identical reticulation sets, and every reticulation is stable for at least one common leaf in both networks, then the in-degrees of the reticulations are guaranteed to be equal in both networks. These conditions ensure that the same system of equations that govern the in-degrees of reticulations holds for both networks and it has a unique solution.
To reiterate, we’re trying to find the sub-class of semi-binary stack-free networks for which equal $\mu$-representations imply equal in-degrees. Now we have shown that this could be a sub-class of reticulation-visible networks, but we need to determine under which conditions the reticulation sets are equal and the reticulations are stable for common leaves. In the case of stack-free orchard networks the reticulations correspond to $\mu$-vectors with multiplicity 2. However, the stability of reticulations with respect to specific leaves may still differ between the networks, even if they share the same $\mu$-representation. Therefore, the next step will be to demonstrate the conditions under which the stability of a node with respect to a leaf $a$ is determined by the $\mu$-representation.
@@ 4.2.1 Preliminary lemmas
Either they are themselves tree-clones, or they are the reticulation parent of a tree-clone, which means by Lemma 9 that they are not stable.
@@
Given a stack-free network $\mathcal{N}$, with $\mu$-representation $\mu(\mathcal{N})$. If $\mu(v)$ has a multiplicity greater than 2 in $\mu(\mathcal{N})$ then all nodes $v\in V$ with $\mu$-vector equal to $\mu(v)$ are not stable.
@@
An edge directed to a reticulation is never a bridge.
@@
There is a path from the root to any parent of a reticulation. A reticulation has at least two parents. If the edge from one of the parents to a reticulation is deleted, then there is still a path from the root to the parent and there is still a path to the reticulation via another parent. Therefore the reticulation and all its descendants are still connected to the root. All other nodes are still connected to the root as well, therefore the graph is still connected. ∎
Let the head of a bridge be called a bridge-node. Then, by the lemma above, bridge-nodes are tree-nodes. All leaves are automatically bridge-nodes, as they become isolated whenever the edge directed into them is deleted. The root is not a bridge-node because it has zero in-degree.
@@
In other words, this lemma implies that a bridge-node is stable for all leaves below it. This means that if $v_{b}$ is a bridge-node, then nodes which lie on a path to a leaf below $v_{b}$ are either ancestors or descendants of $v_{b}$.
@@
Let us first assume that $v_{b}$ is a bridge-node and there is a path $u\rightsquigarrow a$ from an ancestor $u$ of $v_{b}$ to a leaf $a$ below $v_{b}$ which does not pass through $v_{b}$. Then, after deleting the edge from the parent of $v_{b}$ to $v_{b}$, there are still paths from $v_{b}$ to all descendants of $v_{b}$, so all descendants of $v_{b}$ are connected. Furthermore, there is still a path from the root to $a$, which is the product of a path from the root to $u$ and the path $u\rightsquigarrow a$ which does not visit $v_{b}$. Therefore, $a$ is still connected to the root and as all descendants of $v_{b}$ are still connected to each other, this means all descendants are still connected to the root as well. Therefore, the graph is still connected. However, this contradicts our assumption that $v_{b}$ is a bridge-node. Therefore, such a path cannot exist.
Now let us assume all paths from ancestors of $v_{b}$ to leaves below $v_{b}$ pass through $v_{b}$. As $v_{b}$ is a tree-node, there is a single edge directed to $v_{b}$. Therefore, if the edge directed to $v_{b}$ is deleted, there are no longer any paths from ancestors of $v_{b}$ to leaves below $v_{b}$, except those that start in $v_{b}$. The root is an ancestor of $v_{b}$ other than $v_{b}$. Therefore, if the edge directed to $v_{b}$ is deleted, there are no longer any paths from the root to leaves below $v_{b}$, thus the graph is disconnected. This means $v_{b}$ is a bridge-node. ∎
Corollary 3.
Let $v_{b}$ be a bridge-node. Then, for $u\in V$, $u\neq v_{b}$
  • $\mu(u)\geq\mu(v_{b})\iff\mu(u)$ belongs only to ancestors of $v_{b}$
  • $\mu(u)<\mu(v_{b})\iff\mu(u)$ belongs only to descendants of $v_{b}$
Lemma 13.
Let there be a path $u\rightsquigarrow v$ from a node $u$ to a node $v$. If $u\rightsquigarrow v$ is a tree-path, then every path $w\rightsquigarrow v$ ending in $v$ is either contained in $u\rightsquigarrow v$ or contains $u\rightsquigarrow v$.
The proof is given in the appendix.
Lemma 14.
For any node $v$, there is either a tree-path from the root $\rho$ to $v$ or there is a lowest reticulation $r$, such that all paths $\rho\rightsquigarrow v$ pass through $r$.
Proof.
First note, if there is a tree-path from $\rho$ to $v$, then by Lemma 13 there can be no other path $\rho\rightsquigarrow v$ and therefore $v$ has no reticulation ancestors. Now assume there does not exist a tree-path from $\rho$ to $v$. Then $v$ has at least a single reticulation ancestor. Let $r_{l}$ be the lowest reticulation ancestor above $v$, such that $r_{l}$ has no other reticulation descendants which are ancestors of $v$. This $r_{l}$ exists as we only consider finite graphs. Then either $r_{l}$ is equal to $v$, in which case it is clear that all paths from $\rho$ to $v$ visit $r_{l}$, or there is a tree-path from the child $c_{l}$ of $r_{l}$ to $v$. Therefore, by Lemma 13, all paths from ancestors of $c_{l}$ to $v$ must contain $c_{l}\rightsquigarrow v$. Furthermore, $c_{l}$ is a tree-node, therefore $r$ must be its only parent and therefore all paths from ancestors of $r$ to $v$ which contain $c_{l}$ must also contain $r$. The root is an ancestor of $r$. Thus, all paths $\rho\rightsquigarrow v$ pass through $r$. ∎
Lemma 15.
If there is a tree-path from a node $v$ to a bridge-node $v_{b}$, then $v$ is stable with respect to all leaves below $v_{b}$.
Proof.
By Lemma 12, all paths from the root to leaves below $v_{b}$ pass through $v_{b}$. This means that all paths from the root to leaves below $v_{b}$ are a composition of a path from the root to $v_{b}$ and a path from $v_{b}$ to leaf below it. Now let there be a tree-path from a node $v$ to a bridge-node $v_{b}$. Then, by Lemma 13 all paths from ancestors of $v$ to $v_{b}$ pass through $v$. The root is an ancestor of $v$, therefore all paths from the root to $v_{b}$ are a composition of a path from the root to $v$ and the tree-path from $v$ to $v_{b}$. Therefore all paths from the root to leaves below $v_{b}$ are a composition of a path from the root to $v$, the tree-path from $v$ to $v_{b}$ and a path from $v_{b}$ to a leaf below it. In conclusion all paths from the root to leaves below $v_{b}$ pass through $v$, and therefore $v$ is stable with respect to all leaves below $v_{b}$. ∎
Lemma 16.
Proof.
Let us assume there is a tree-path from the child $c_{r}$ of a reticulation $r$ to a bridge-node $v_{b}$. By Lemma 15, $c_{r}$ is stable with respect to all leaves below $v_{b}$. Then, by Lemma 9, $r$ is stable with respect to all leaves below $v_{b}$. ∎
Let us define the set $A_{b}$ as the subset of the ancestors of a bridge-node $v_{b}$, whose $\mu$-vectors have multiplicity 2 in $\mu(\mathcal{N})$. For any set $S\subseteq V$, let $\mu(S)$ be the multiset of $\mu$-vectors of the nodes in $S$. Then $\mu(A_{b})$ is exactly the multiset $\{\mu(v):\mu(v)\geq\mu(v_{b}),\#\mu(v)=2\}$ and by Corollary 3, each vector in $\mu(A_{b})$ only belongs to ancestors of $v_{b}$. Notice that $\mu(A_{b})$ is purely defined in terms of $\mu$-vectors if given a $\mu(v_{b})$ which belongs to a bridge-node, and it can be determined only from the $\mu$-representation, without any knowledge of the graph. We will show that it can be used to find $\mu$-vectors which belong to stable reticulations.
Lemma 17.
If $\mu(v)$ minimal in $\mu(A_{b})$, then it belongs to a reticulation and to the child of that reticulation.
Proof.
A $\mu$-vector with multiplicity 2 either belongs to a pair of tree-clones or to a reticulation and its child. Now, if $\mu(v_{1})\in\mu(A_{b})$ belongs to a pair of tree-clones $v_{1},v_{2}$ with $\mu(v_{1})=\mu(v_{2})$, then there cannot be a tree-path from either of them to a bridge-node $v_{b}$, because by Lemma 8, tree-clones are not stable. Therefore, $v_{1}$ and $v_{2}$ must have reticulation descendants $r_{1},r_{2}$, who are ancestors of $v_{b}$. Note that $r_{1}$ could be equal to $r_{2}$. If the multiplicity of either $r_{1}$ or $r_{2}$ is greater than 2, then their child must be a tree-clone, by the contrapositive of Lemma 10. By the same argument there must then be other reticulation descendants of $v_{1}$ and $v_{2}$ above $v_{b}$. We only consider finite graphs, and therefore, w.l.o.g. we can assume $\mu(r_{1})$ and $\mu(r_{2})$ have multiplicity 2 in $\mu(\mathcal{N})$. Then, $r_{1},r_{2}\in A_{b}$, with $\mu(r_{1})<\mu(v_{1})$ and $\mu(r_{2})<\mu(v_{2})$. Therefore, $\mu(v_{1})$ is not minimal in $A_{b}$. We can conclude that if $\mu(r)$ is minimal in $A_{b}$, then $r$ is a reticulation. ∎
Lemma 18.
There is a tree-path from the child of a reticulation $r$ to a bridge-node $v_{b}$ if, and only if, $\mu(r)$ is minimal in the set $\mu(A_{b})$.
Proof.
@@
For the first direction, assume there is a tree-path from the child of a reticulation $r$ to a bridge $v_{b}$. Then, by Lemma 16, $r$ is stable. Therefore, by Corollary 2, $\mu(r)$ has multiplicity 2 in $\mu(\mathcal{N})$. Furthermore, $r$ is an ancestor of $v_{b}$, therefore $r\in A_{b}$. Also note, that by Lemma 13 and Lemma 14, $r$ is the lowest reticulation above $v_{b}$. Then, by combining Lemma 14 and Lemma 17, all ancestors of $v_{b}$, with $\mu$-vectors which have multiplicity 2, are ancestors of $r$. Thus, we have $\mu(v)\geq\mu(r)$, for $\mu(v)\in\mu(A_{b})$. Therefore, $\mu(r)$ is minimal in $A_{b}$. This proves the first direction.
@@
Now assume $\mu(r)$ is minimal in $\mu(A_{b})$. By Lemma 17, $\mu(r)$ belongs to a reticulation $r$ and its child $c_{r}$. Furthermore, because $\mu(c_{r})=\mu(r)\in\mu(A_{b})$, both $r$ and $c_{r}$ are ancestors of $v_{b}$. Therefore, there is a path from $c_{r}$ to $v_{b}$. If the path from $c_{r}$ to $v_{b}$ is not a tree-path, then $r$ is not the lowest reticulation above $v_{b}$. Let $r_{\ell}$ be the lowest reticulation above $v_{b}$, then by Lemma 14, $r_{\ell}$ is a descendant of $r$. By Lemma 18, $r_{\ell}$ is stable. Therefore, by Corollary 2, #$\#\mu(r_{l})\leq 2$. The child of $r_{\ell}$ has the same $\mu$-vector, therefore #$\#\mu(r_{l})=2$. This means that $\mu(r_{\ell})\in\mu(A_{b})$ and because $r_{\ell}$ is a descendant of $r$, also $\mu(r)>\mu(r_{\ell})$. This contradicts our assumption that $\mu(r)$ is minimal in $\mu(A_{b})$. Therefore, $r$ is the lowest reticulation above $v_{b}$, and the path from $c_{r}$ to $v_{b}$ is a tree-path. ∎
@@ 4.2.2 Strongly reticulation-visible networks
The lemmas in the previous section show that if we know that a given $\mu$-vector belongs to a bridge-node, then we can find the $\mu$-vector of the lowest reticulation above this bridge-node and the reticulation will be stable for all leaves below the bridge-node. We therefore propose to consider the class of networks such that for each reticulation there is a tree-path from its child to a bridge. We will call this the class of strongly reticulation-visible networks.
@@
Note that all strongly reticulation-visible networks are reticulation-visible. But there are reticulation-visible networks which are not strongly reticulation-visible. See Figure 10 for an example. Therefore the class of strongly reticulation-visible networks is a proper subclass of the class of reticulation-visible networks, which is itself a proper subclass of stack-free networks. If we can show that for strongly reticulation-visible networks it is possible to determine whether a $\mu$-vector belongs to a bridge-node from just the $\mu$-representation (Proposition 3), we will have shown that the stability of reticulations can be determined. Knowing the stability of reticulations will show that the reticulation in-degrees can be determined (Proposition 2), finally leading to our main isomorphism result (Theorem 4).

### v2

@@ 4.2 Determining stable nodes
In this subsection, we work with stack-free phylogenetic networks, denoted by $\mathcal{N}=(V,E)$. 2 shows that if two reticulation-visible networks have equal $\mu$-representations, identical reticulation sets, and every reticulation is stable for at least one common leaf in both networks, then the in-degrees of the reticulations are guaranteed to be equal in both networks. These conditions ensure that the same system of equations that govern the in-degrees of reticulations holds for both networks and it has a unique solution.
To reiterate, we’re trying to find the sub-class of semi-binary stack-free networks for which equal $\mu$-representations imply equal in-degrees. Now we have shown that this could be a sub-class of reticulation-visible networks, but we need to determine under which conditions the reticulation sets are equal and the reticulations are stable for common leaves. In the case of stack-free orchard networks the reticulations correspond to $\mu$-vectors with multiplicity 2. However, the stability of reticulations with respect to specific leaves may still differ between the networks, even if they share the same $\mu$-representation. Therefore, the next step will be to demonstrate the conditions under which the stability of a node with respect to a leaf is determined by the $\mu$-representation.
@@ 4.2.1 Preliminary lemmas
Either they are themselves tree-clones, or they are the reticulation parent of a tree-clone, which means by 9 that they are not stable.
@@
Let $\mathcal{N}$ be a stack-free network with $\mu$-representation $\mu(\mathcal{N})$. If $\mu(v)$ has a multiplicity greater than 2 in $\mu(\mathcal{N})$ then all nodes $v\in V$ with $\mu$-vector equal to $\mu(v)$ are not stable.
@@
Let the head of a bridge be called a bridge-node. Then, since reticulation edges are never bridges, bridge-nodes are tree-nodes. All leaves are automatically bridge-nodes, as they become isolated whenever the edge directed into them is deleted. The root is not a bridge-node because it has zero in-degree.
Observation 2.
In other words, this observation implies that a bridge-node is stable for all leaves below it. This means that if $v_{b}$ is a bridge-node, then nodes which lie on a path to a leaf below $v_{b}$ are either ancestors or descendants of $v_{b}$.
Observation 3.
Let $v_{b}$ be a bridge-node. Then, for $u\in V$, $u\neq v_{b}$
  • $\mu(u)\geq\mu(v_{b})\iff\mu(u)$ belongs to an ancestor of $v_{b}$.
  • $\mu(u)<\mu(v_{b})\iff\mu(u)$ belongs to a descendant of $v_{b}$.
Observation 4.
Let us define the set $A_{b}$ as the subset of the ancestors of a bridge-node $v_{b}$, whose $\mu$-vectors have multiplicity exactly 2 in $\mu(\mathcal{N})$. While we have defined $\mu$ on networks and on individual nodes, we generalize this definition to a set of vertices. For any set $S\subseteq V$, let $\mu(\mathcal{N})_{S}$ be the multiset of $\mu$-vectors of the nodes in $S$. In other words, $\mu(\mathcal{N})_{S}=\{\mu(v)\in\mu(\mathcal{N}):v\in S\}$. Then $\mu(\mathcal{N})_{A_{b}}$ is exactly the multiset $\{\mu(v):\mu(v)\geq\mu(v_{b}),\#\mu(v)=2\}$. By 3, each vector in $\mu(\mathcal{N})_{A_{b}}$ only belongs to ancestors of $v_{b}$. Notice that $\mu(\mathcal{N})_{A_{b}}$ is purely defined in terms of $\mu$-vectors if given a $\mu(v_{b})$ which belongs to a bridge-node, and it can be determined from the $\mu$-representation, without any knowledge of the graph. Here, we show that it can be used to find $\mu$-vectors which belong to stable reticulations. In Section 4.2.2, we show how to determine $\mu$-vectors that belong to bridge-nodes.
@@
If $\mu(v)$ is minimal in $\mu(\mathcal{N})_{A_{b}}$, then it belongs to a reticulation and to the child of that reticulation.
@@
A $\mu$-vector with multiplicity 2 either belongs to a pair of tree-clones or to a reticulation and its child. Now, if $\mu(v_{1})\in\mu(\mathcal{N})_{A_{b}}$ belongs to a pair of tree-clones $v_{1},v_{2}$ with $\mu(v_{1})=\mu(v_{2})$, then there cannot be a tree-path from either of them to a bridge-node $v_{b}$, because by 8, tree-clones are not stable. Therefore, $v_{1}$ and $v_{2}$ must have reticulation descendants $r_{1},r_{2}$, who are ancestors of $v_{b}$. Note that $r_{1}$ could be equal to $r_{2}$. If the multiplicity of either $r_{1}$ or $r_{2}$ is greater than 2, then their child must be a tree-clone, by the contrapositive of 10. By the same argument there must then be other reticulation descendants of $v_{1}$ and $v_{2}$ above $v_{b}$. We only consider finite graphs, and therefore, w.l.o.g. we can assume $\mu(r_{1})$ and $\mu(r_{2})$ have multiplicity 2 in $\mu(\mathcal{N})$. Then, $r_{1},r_{2}\in A_{b}$, with $\mu(r_{1})<\mu(v_{1})$ and $\mu(r_{2})<\mu(v_{2})$. Therefore, $\mu(v_{1})$ is not minimal in $A_{b}$. We can conclude that if $\mu(r)$ is minimal in $A_{b}$, then $r$ is a reticulation. ∎
@@
There is a tree-path from the child of a reticulation $r$ to a bridge-node $v_{b}$ if, and only if, $\mu(r)$ is minimal in the set $\mu(\mathcal{N})_{A_{b}}$.
@@
For the first direction, assume there is a tree-path from the child of a reticulation $r$ to a bridge $v_{b}$. Then, by 4, $r$ is stable. Therefore, by 2, $\mu(r)$ has multiplicity 2 in $\mu(\mathcal{N})$. Furthermore, $r$ is an ancestor of $v_{b}$, therefore $r\in A_{b}$. Also note, that $r$ is the lowest reticulation above $v_{b}$. Then, by 11, all ancestors of $v_{b}$, with $\mu$-vectors which have multiplicity 2, are ancestors of $r$. Thus, we have $\mu(v)\geq\mu(r)$, for $\mu(v)\in\mu(\mathcal{N})_{A_{b}}$. Therefore, $\mu(r)$ is minimal in $A_{b}$. This proves the first direction.
@@
The other direction we will prove by contradiction. Assume $\mu(r)$ is minimal in $\mu(\mathcal{N})_{A_{b}}$. By 11, $\mu(r)$ belongs to a reticulation $r$ and its child $c_{r}$. Furthermore, because $\mu(c_{r})=\mu(r)\in\mu(\mathcal{N})_{A_{b}}$, both $r$ and $c_{r}$ are ancestors of $v_{b}$. Therefore, there is a path from $c_{r}$ to $v_{b}$. If the path from $c_{r}$ to $v_{b}$ is not a tree-path, then $r$ is not the lowest reticulation above $v_{b}$. Let $r_{\ell}$ be the lowest reticulation above $v_{b}$, then $r_{\ell}$ is a descendant of $r$. By 4, $r_{\ell}$ is stable. Therefore, by 2, #$\#\mu(r_{l})\leq 2$. The child of $r_{\ell}$ has the same $\mu$-vector, therefore #$\#\mu(r_{l})=2$. This means that $\mu(r_{\ell})\in\mu(\mathcal{N})_{A_{b}}$ and because $r_{\ell}$ is a descendant of $r$, also $\mu(r)>\mu(r_{\ell})$. This contradicts our assumption that $\mu(r)$ is minimal in $\mu(\mathcal{N})_{A_{b}}$. Therefore, $r$ is the lowest reticulation above $v_{b}$, and the path from $c_{r}$ to $v_{b}$ is a tree-path. ∎
Lemma 12 shows that given a vector $\mu(v_{b})$ which belongs to a bridge-node, we can determine the $\mu$-vector belonging to the lowest reticulation above that bridge by finding the minimal $\mu$-vector $\mu(v)\geq\mu(v_{b})$ with multiplicity 2.
@@ 4.2.2 Strongly reticulation-visible networks
The lemmas in the previous section show that if we know that a given $\mu$-vector belongs to a bridge-node, then we can find the $\mu$-vector of the lowest reticulation above this bridge-node and the reticulation will be stable for all leaves below the bridge-node (which are the leaves for which the $\mu$-vector of the bridge-node has non-zero coordinates). We therefore propose to consider the class of networks such that for each reticulation there is a tree-path from its child to a bridge. We will call this the class of strongly reticulation-visible networks.
@@
Note that all strongly reticulation-visible networks are reticulation-visible. But there are reticulation-visible networks which are not strongly reticulation-visible. See Figure 10 for an example. Therefore the class of strongly reticulation-visible networks is a proper subclass of the class of reticulation-visible networks, which is itself a proper subclass of stack-free networks. If we can show that for strongly reticulation-visible networks it is possible to determine whether a $\mu$-vector belongs to a bridge-node from just the $\mu$-representation (3), we will have shown that the stability of reticulations can be determined. Knowing the stability of reticulations will show that the reticulation in-degrees can be determined (2), finally leading to our main isomorphism result (Theorem 4).